Problem 1

Var rate;

Var hours;

If (hours >= 40)

               Print ((rate * 40) + rate);

Else

               Print (rate * 40);

What is the answer?

1.      If rate = 850

Hours = 40

Answer =

2.      If rate = 850

Hours = 20

Answer =


1
Expert's answer
2021-10-18T18:28:22-0400

1..

Answer=(rate * 40) + rate)

=(850*40)+850

=34850

2..

Answer=(rate * 40)

=850*40

=34000